OpenStudy (anonymous):

need help with 54-6divided by 2+6=?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

it equals 6 i think?

OpenStudy (anonymous):

no the answe is 57 i dont know how the got it 54-6 div by 2 plus 6 =

OpenStudy (anonymous):

it 30

OpenStudy (anonymous):

57

OpenStudy (bahrom7893):

54 - (6/2) + 6 = 54 - 3 + 6 = 60 - 3 = 57

OpenStudy (bahrom7893):

Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)